OpenStudy (anonymous):

what is the equation in standard form? x^2+y^2-12x-12y+36=0

OpenStudy (anonymous):

do you know the equation for standard form

OpenStudy (anonymous):

idk if its an ellipse or what

OpenStudy (anonymous):

its a circle and the equation would be (x - h)² + (y - k)² = r²

OpenStudy (anonymous):

you can complete the square to write it in standard form do you know how to complete the square

OpenStudy (anonymous):

i do not waaaaa

OpenStudy (anonymous):

okay Take half of 12, the coefficient of x, which is 6. Then square 6 and get 36. Now add 36 to both sides: x² + 12x + 36 + y² + 12y = -36 + 36 Complete the square on the y's: Take half of 12, the coefficient of y, which is 6. Then square 6 and get 36. Now add 36 to both sides: x² + 12x + 36 + y² + 12y + 36 = -36 + 36 + 36 Factor the first three terms as (x + 6)(x + 6) and as (x + 6)² and combine the terms on the right (x + 6)² + y² + 12y + 36 = 36 Factor the next three terms as (y + 6)(y + 6) and as (y + 6)² (x + 6)² + (y + 6)² = 36 Now we compare with the standard form: (x - h)² + (y - k)² = r² And see that -h = +6 or h = -6, -k = 6, so k = -6, and r² = 36 so r = 6.

OpenStudy (anonymous):

why thank you very much

OpenStudy (anonymous):

your welcome:)